Herbert and Dranreb can thoroughly clean a house in 3 hours. If each work, one of them would do the job in 1 hour less time than the other. How long would it take each person to do the cleaning?
1
Expert's answer
2020-05-12T16:58:26-0400

First of all let's write collaboration time formula:

"t_C=\dfrac{V}{\upsilon_1+\upsilon_2}" , where tC - collaboration time, V - volume of the work, "\upsilon_1" - labor productivity of the first man, "\upsilon_2" - labor productivity of the second man.

Now let's write labor productivity formulas for both men:

"\upsilon_1=\dfrac{V}{t_1}" and "\upsilon_2=\dfrac{V}{t_2}" , where V - volume of the work, t1 - first man-hours used, t2 - second man-hours used.

Let's put these formulas in collaboration time formula:

"t_C=\dfrac{V}{\dfrac{V}{t_1}+\dfrac{V}{t_2}}=\dfrac{V}{V(\dfrac{1}{t_1}+\dfrac{1}{t_2})}="


"=\dfrac{1}{\dfrac{1}{t_1}+\dfrac{1}{t_2}}=\dfrac{1}{\dfrac{t_2+t_1}{t_1*t_2}}=\dfrac{t_1*t_2}{t_1+t_2}"


Let's mark t1 as x, then t2 will be x - 1. Collaboration time is equal to 3 hours. Now we can write equation:

"\dfrac{x*(x-1)}{x+(x-1)}=3"

Now let's simplify and solve it:

"\dfrac{x^2-x}{2x-1}=3"


"x^2-x=3(2x-1)"

"x^2-x=6x-3"

"x^2-7x+3=0"

a = 1, b = -7, c = 3

"x_1=\dfrac{-b-\sqrt{b^2-4ac}}{2a}="


="\dfrac{-(-7)-\sqrt{(-7)^2-4*1*3}}{2*1}=\dfrac{7-\sqrt{49-12}}{2}="

"=\dfrac{7-\sqrt{37}}{2}\approx0.5" hours does not satisfy the conditions of our task because man alone can't clean house faster than both men together (0.5 < 3)

"x_2=\dfrac{-b+\sqrt{b^2-4ac}}{2a}="


="\dfrac{-(-7)+\sqrt{(-7)^2-4*1*3}}{2*1}=\dfrac{7+\sqrt{49-12}}{2}="


"=\dfrac{7+\sqrt{37}}{2}\approx6.5" hours is man-hours of the first guy.


6.5 - 1 = 5.5 hours is man-hours of the second guy.


Answer: 6.5 hours and 5.5 hours or 6 hours 30 mins and 5 hours 30 mins.